OpenStudy (hapahearthero):

Who can help with piecewise functions? Question is about electricity rates charges. Base charge is $7.00 First 400kWh or less at 0.0450/kWh Over 400 kWh at 0.0750/kWh I know that w(x)=7.00+0.0450x if 400 or less kWh I am confused how to write out the solution for over 400kWh. The solution is telling me it is: -5+0.0750x but where did the -5 come from?

jimthompson5910 (jim_thompson5910):

use w(x)=7.00+0.0450x to find the value of w(x) when x = 400 what do you get?

OpenStudy (hapahearthero):

Yes I understand how to form the functions for up to 400 kWh but it's the over 400 kWh part that is tricking me out

OpenStudy (hapahearthero):

I figured it out! the solution to 7.00+0.045x when x is 400 is 25. Then we would add the 25 to the next values =25+0.075(x-400) ..... then distribute whats outside the parenthesis to whats inside =25+0.075x-30...... then combine like terms(in this case just adding 25+ -30 to get -5 =-5+0.075x Yay

jimthompson5910 (jim_thompson5910):

Or you can think of it like this some linear function y = mx+b goes through (400,25). We know the slope for that piece is 0.075, so m = 0.075 plug in m = 0.075, x = 400, y = 25. Then solve for b y = mx+b 25 = 0.075*400+b 25 = 30+b 25-30 = 30+b-30 -5 = b b = -5 So y = mx+b turns into y = 0.075x-5 which is the same as y = -5+0.075x

OpenStudy (hapahearthero):

Oh man amazing. That is a great way to find it also, I am going to write this down for the future problems. Thank you!!!!

jimthompson5910 (jim_thompson5910):

you're welcome